OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. 3-n/2=3/4

OpenStudy (anonymous):

subtract 3: n/2=3/4-3 rewrite: n/2 = 3/4 -12/4 simplify: n/2 = -9/4 times by 2: n=-18/4 simplify: n=-9/2

OpenStudy (anonymous):

can u simplify -9/2?

OpenStudy (anonymous):

-4.5

OpenStudy (anonymous):

otherwise no.